# Normality in the square of the Sorgenfrey line

### v1

@@ Abstract.
We consider sets of reals $X$ endowed with the Sorgenfrey lower limit topology denoted $X[\leq]$. Przymusiński proved that if $X$ is a $Q$-set then $(X[\leq])^{2}$ is normal. While the converse is not in general true we consider examples of sets of the reals for which $(X[\leq])^{2}$ is normal or just pseudo-normal. For example, if $X$ is a $\lambda$ set, then $(X[\leq])^{2}$ is pseudo-normal but assuming CH there is an $X$ concentrated on a countable dense subset (so not a $\lambda$-set) but still $(X[\leq])^{2}$ is normal.
@@
$Q$-set, $\lambda$-set, Sorgenfrey topology
@@
Primary:
@@ 1. Introduction
A $Q$-set is a subset of the reals with the property that every subset is $G_{\delta}$ (equivalently $F_{\sigma}$) in the relative subspace topology. The notion was first introduced in [hausdorff1933] and $Q$-sets were used to produce consistent counterexamples to the Normal Moore Space Conjecture [Bing1951]. For example, if $X\subseteq 2^{\omega}$, the set of branches $b_{x}=\{x\upharpoonright n\}$ in $2^{<\omega}$ is an almost disjoint family and the corresponding $\Psi$-space on $2^{<\omega}$ is normal if and only if $X$ is a $Q$-set. Similarly, if $X$ is a subset of the $x$-axis in the upper half plan of ${\mathbb{R}}^{2}$, then the Moore-Niemitzky Plane topology (aka the bubble space) is normal if and only if $X$ is a Q-set. For more on these examples see [T]. On the other hand Prymusiński [T.Przymusinski1973] proved that if $X$ is a Q-set endowed with the Sorgenfrey Topology then $X^{2}$ is normal.
@@
We first observe that the converse to Przymusiński’s theorem does not hold: Todorčević observed that if $X$ is an entangled set of reals and is endowed with the Sorgenfrey topology, then every finite power is both hereditarily Lindelöf and hereditarily separable [Todorcevic1989]. Recall that CH implies both the existence of entangled sets of reals and the non-existence of $Q$-set.
There is the weaker notion of a set of reals $X$ being a $\lambda$-set. $X$ is a $\lambda$-set if every countable subset is $G_{\delta}$. While the existence of a $Q$-set of reals is independent, there are $\lambda$-sets in ZFC. Analagous to the $Q$-set results mentioned above, the $\Psi$-space over the set of branches of a subset $X\subseteq 2^{\omega}$ is pseudo-normal (meaning that every pair of closed sets, one of which is countable, can be separated by disjoint open sets) if and only if $X$ is a $\lambda$-set and also $X$ is a $\lambda$-set if and only if the Moore-Niemytzki plane over $X$ is pseudo-normal (see [vD])
We provide some partial converses to Przymusiński’s theorem that underscore what is needed to prove establish normality of $(X[\leq])^{2}$. We show that the expected analogous result for the Sorgenfrey topology on a $\lambda$-set holds. Namely, if $X$ is a $\lambda$-set of reals, then $(X[\leq])^{2}$ is pseudonormal. We also construct, assuming CH, a set of reals for which $X([\leq])^{2}$ is normal such that $X$ is concentrated on a countable dense subset, and so not a $\lambda$-set.
@@
Let $X$ be a topological space. Let $A,B$ be subsets in $X$, AA has the covering property with respect to $B$ iff there exists a countable cover $\{U_{n}\}_{n<\omega}$ such that $\overline{U_{n}}\cap B=\emptyset$. If there is no ambiguity in the context, we just say $A$ has the covering property.
@@
All of our topological terminology is standard and we refer the reader to Engelking for any undefined notions.
@@ 2. Q-sets and the Sorgenfrey square
Let $A,B$ be subsets of reals. For any strictly decreasing function $f:A\to B$, there is a countable $A_{0}$ such that $f\upharpoonright(A\setminus A_{0}):A\setminus A_{0}\rightarrow f(A\setminus A_{0})$ is a homeomorphism.
@@
Suppose $X$ is a set of reals. Suppose further that
@@
    there exists a strictly decreasing homeomorphism $f:X\to X$, and
@@
If there exists $\mathcal{U},\mathcal{V}\subseteq X^{2}$ Sorgenfrey open such that $\mathcal{U}\supseteq G_{f|_{A}},\mathcal{V}\supseteq G_{f_{B}}$ and $\mathcal{U}\cap\mathcal{V}=\emptyset$, then $A$ is relatively $F_{\sigma}$ in $X$.
@@
Note that $A=\cup_{n<\omega}A_{n}$ where $A_{n}=\{a\in A:U((a,f(a)),1/n)\subseteq\mathcal{U}\}$, we claim that
@@
$\overline{A_{n}}^{E}\cap X\subseteq A$
Proof.
Assume otherwise, if $a_{n}\to b$ for $a_{n}\in A$ and $b\in B$. There exists an $r>0$ such that $U((b,f(b)),r)\subseteq\mathcal{V}$. By continuity of $f$, we have $f(a_{n})\to f(b)$ as well. Let $\epsilon=\min(1/n,r)$, Wlog there exists $a_{n_{0}}$ such that $b<a_{n_{0}}<b+\epsilon$ and $f(b)<f(a_{n_{0}})<f(b)+\epsilon$. Since $\epsilon<1/n$, $(a_{n_{0}},f(a_{n_{0}})+\epsilon)$ is a point in $U((a_{n_{0}},f(a_{n_{0}})),1/n)\subseteq\mathcal{U}$ However, we have
$f(b)<f(a_{n_{0}})+\epsilon<f(b)+2\epsilon<f(b)+r$
So $(a_{n_{0}},f(a_{n_{0}}))$ is also a point in $U((b,f(b)),r)\subseteq\mathcal{V}$, but $\mathcal{U}$ and $\mathcal{V}$ should have empty intersection hence this is a contradiction. ∎
Now with this claim, we have $A=\cup_{n<\omega}(\overline{A_{n}}^{E}\cap X)$ which is a relatively $F_{\sigma}$ set. ∎
@@
Let $(X[\leq])^{2}$ be normal. If there exists a strictly decreasing homeomorphism $f:X\to X$, then $X$ is a $Q$-set.
@@
Since $f$ is a homeomorphism, its graph $G_{f}$ is closed and discrete. Hence any two subsets of $G_{f}$ is closed. In particular, take any $A\subseteq X$, $G_{f|_{A}}$ and $G_{f|_{X\setminus A}}$ can be separated by open sets. By 2.2, $A$ is relatively $F_{\sigma}$ in $X$. Hence $X$ is a $Q$-set. ∎
@@
By 2.1, there exists a countable set $C$ such that $f|_{X\setminus C}:X\setminus C\to X\setminus C$ is a homeomorphism. Then let $g=f|_{X\setminus C}$, the graph $G_{g}$ of $g$ is closed and discrete. Since $(X[\leq])^{2}$ is normal, any two subsets of $G_{g}$ can be separated by Sorgenfrey open subsets of $X^{2}$. In particular, let $X^{\prime}=X\setminus C$, take any subset $A\subseteq X^{\prime}$, $G_{g|_{A}}$ and $G_{g|_{X\prime\setminus A}}$ can be separated by open sets $\mathcal{U},\mathcal{V}$. Take their intersection to $(X^{\prime})^{2}$, $\mathcal{U}\cup(X^{\prime})^{2}$ and $\mathcal{V}\cup(X^{\prime})^{2}$ are still Sorgenfrey open in $(X^{\prime})^{2}$. By 2.2, $A$ is relatively $F_{\sigma}$ in $X^{\prime}$, hence $X^{\prime}$ is a $Q$ set, and it only lacks countably many points. ∎
Baumgartner proved that PFA implies that any two $\aleph_{1}$-dense sets of reals are order isomorphic [Baumgartner1973]. This consequence of PFA has come to be referred to as Baumgartner’s Axiom (BA). From the proof of the above proposition we obtain the following corollary:
@@
Any uncountable $X$ contains an $\aleph_{1}$-dense subset. By BA it is order isomorphic to its reverse order. So by the proof of Proposition 2.5 there is a countable $C$ such that $X\setminus C$ must be a $Q$-set.∎
@@
Let $f:A\to\mathbb{R}$ be strictly decreasing on uncountable $A\subseteq\mathbb{R}$, then $f$ can be extended to strictly decreasing function on $\overline{A}\setminus C$ for some countable set $C$.
@@
$f$ can be extended to a decreasing function on $\overline{A}$. Note that this function is not necessarily strictly decreasing. By 2.1, we can remove a countable set $B_{0}$ such that $f$ is homeomorphism on $\overline{A}\setminus f^{-1}(B_{0})$ hence strictly decreasing. We will show $f^{-1}(B_{0})$ is countable by showing for every $b\in B_{0}$, $f^{-1}(b)$ has at most 3 points. Assume there are 4 points $x<y<z<w$ in $f^{-1}(b)$. Note that between $x$ and $w$ there is at most 1 point in $A$, since for every $z_{1}<z_{2}\in A$,
@@
but $f$ is strictly decreasing on $A$, so there can only be at most 1 point in $A$. Hence, there are no points in $\overline{A}\setminus A$. But $y,z$ can not be both in $A$, we have a contradiction. Let $C=f^{-1}(B_{0})$, we have the result. ∎
Recall a set $L\subseteq\mathbb{R}$ is $\mathbf{n}$-entangled iff for any set of disjoint $n$-tuples $A\subseteq L^{n}$ of size $\omega_{1}$ and for every $s\in 2^{n}$ there exists $x,y\in A$ such that $\forall i\leq n(x_{i}<y_{i})\iff s_{i}=0$. $L$ is entangled iff $L$ is $\mathbf{n}$-entangled for all $n<\omega$. Let $s\in 2^{n}$, $A\subseteq L^{n}$ is said to have type $s$ iff $\forall i\leq n(x_{i}<y_{i})\iff s_{i}=0$.
The following lemma is due to [Todorcevic1985].
@@
[Todorcevic1985] Let $E\subseteq\mathbb{R}$ to have the following property:
  • No one-to-one continuous function can be defined from $A\to E$ for any uncountable $A\subseteq E^{n}$, for any $n<\omega$
@@
Proof.
Assume $E$ is not entangled, let $m<\omega$ be the least such that $E$ is $m$-entangled, but not $m+1$-entangled. Then there exists an uncountable $A\subseteq E^{m+1}$ such that $A$ is missing one type $s\in 2^{n}$, with $s=(s_{1},s_{2},\dots,s_{m},s_{m+1})$. Wlog, let $s_{m+1}=0$. Define $f:E^{m}\to E$ by $f(z)=r$ iff $z\text{\textasciicircum}r\in A$. We will show $f$ is continuous on some uncountable subset of $dom(f)$, which is not possible by the property $E$ has. For any $p\in dom(f)$, define
$\omega(p)=\cap\{\overline{f(dom(f)\cap I)}:I\text{ open in }E^{m},p\in I\}$
$\omega(p)$ is all the possible limit at the point $p$, if $|\omega(p)|=0$ or $1$, $f$ is continuous at $p$. Let $B=\{p\in dom(f):|\omega(p)|\geq 2\}$, assume $B$ is uncountable. By countable chain condition and symmetry, there exists an uncountable subset $B_{0}$ such that
$\forall z\in B_{0},\exists r\in\omega(z)\;f(z)<d<r$
Note that $B_{0}\subseteq E^{m}$, and since $E$ is $m$-entangled, there exists $u,v\in B_{0}$, $u=(u_{1},\dots,u_{m}),v=(v_{1},\dots,v_{m})$, such that $\forall i\leq m(u_{i}<v_{i})\iff s_{i}=0$. Let
$I=\{z\in E^{m}:\forall i\leq m(u_{i}<z_{i})\iff s_{i}=0\}$
II is an open set. However for all $z\in dom(f)\cap I$, since $s_{m+1}=0$ and $A$ is missing this type, we can only have $f(z)<f(u)<d$, and $\omega(v)\subseteq\overline{f(dom(f)\cap I)}\implies\omega(v)<d$ which is a contradiction. ∎
@@
    No one-to-one continuous function can be defined from $A\to E$ for any uncountable $A\subseteq E^{n}$, for any $n<\omega$.
@@
    $E$ contains $\mathbb{Q}$ and $\mathbb{Q}$ is not a relative $G_{\delta}$-set in $E$. Moreover, $E$ can be chosen so that it is concentrated on $\mathbb{Q}$. I.e., $E\setminus U$ is countable for every open set $U\supseteq\mathbb{Q}$.
@@
$\mathcal{F}=\{$ $f:U\to\mathbb{R}:f\text{ one-to-one continuous function on Borel set }U\subseteq\mathbb{R}^{n},n<\omega,$
$\text{and let }U=\Pi_{i=1}^{n}U_{i},\text{we have }\cup_{i=1}^{n}U_{i}\text{ uncountable}\}$
@@
For every $\alpha<\omega_{1}$ define the following set:
$C_{\alpha}$ $=\bigcap_{\beta\leq\alpha}A_{\alpha}\setminus\mathbb{Q}.$
Note that $C_{\alpha}$ is uncountable by the Baire Category Theorem. We now define elements $x_{\alpha}\in\mathbb{R}$ recursively as follows:
Stage 0: Take $x_{0}\in C_{0}$ such that $f_{0}(\{x_{0}\}^{n})\not=x_{0}$
$X_{\alpha}$ $=\{x_{\beta}:\beta<\alpha\}$
$I_{\alpha}$ $=\{f_{\beta}(x_{\gamma_{1}},x_{\gamma_{2}},\dots,x_{\gamma_{n}}):\beta\leq\alpha,\gamma_{i}<\alpha,i\leq n\}$
$P_{\alpha}$ $=\cup\{f_{\beta}^{-1}(x_{\gamma}):\beta\leq\alpha,\gamma<\alpha\}$
Note that all of the above sets are countable, and $C_{\alpha}$ is uncountable. Hence, we can choose $x_{\alpha}\in C_{\alpha}\setminus(X_{\alpha}\cup I_{\alpha}\cup P_{\alpha})$, such that either $x_{\alpha}\not\in\cup dom(f_{\alpha})$ or $f_{\alpha}([X_{\alpha}]^{n}\cap dom(f))\not\ni x_{\alpha}$.
@@
Verify Property 1: Fix $n<\omega$, for any one to one continuous function $g:A\to E$ defined on some uncountable set $A\subseteq E^{n}$, $g$ can always be extended to a continuous function on a Borel set. , i.e., there is an $\alpha_{0}$ such that $f_{\alpha_{0}}\upharpoonright A=g$. However, by choice of $x_{\alpha}$, for all $\alpha>\alpha_{0}$, we have that $f_{\alpha_{0}}(\{x_{\gamma}:\gamma<\alpha\}^{n})$ can not take on any values in $\{x_{\beta}:\beta>\alpha_{0}\}$. Indeed for all $\beta>\alpha_{0}$, if $\alpha_{0}<\beta<\alpha$, we have $x_{\alpha}\not\in P_{\alpha}\implies x_{\alpha}\not\in\cup f^{-1}(x_{\beta})\implies f_{\alpha_{0}}(\{x_{\gamma}:\gamma<\alpha\}^{n})\neq x_{\beta}$. If $\alpha_{0}<\alpha<\beta$, we have $x_{\beta}\not\in I_{\beta}\implies x_{\beta}\not\in f(\{x_{\gamma}:\gamma<\beta\}^{n})\implies x_{\beta}\not\in f(\{x_{\gamma}:\gamma<\alpha\}^{n})$. So for all $z\in E^{n}$ such that $z_{i}=x_{\alpha}$ for some $i\leq n$, $f_{\alpha_{0}}(z)\in\mathbb{Q}\cup\{x_{\beta}:\beta\leq\alpha_{0}\}$ which is a countable set. Hence $g^{-1}(E)$ is countable contradicting the fact that $g$ was defined on the uncountable set $A$.
Verify Property 2: That $X$ is concentrated on $\mathbb{Q}$ follows easily from the fact that for every $\alpha$, $X\setminus A_{\alpha}\subseteq\{x_{\beta}:\beta<\alpha\}$. In particular, $\mathbb{Q}$ is not $G_{\delta}$ in $X$ and so $X$ is not a $\lambda$-set.
By 2.7, the set $E$ we constructed in 2.8 is entangled and in particular $2$-entangled. By 2.6 this implies no monotone functions can be defined on $E$. Hence we have the following result.

### v2

# Submitted toTopology Proceedings

@@ Abstract.
We consider sets of reals X endowed with the Sorgenfrey lower limit topology denoted $X[\leq]$. Przymusiński proved that if X is a Q-set then $(X[\leq])^{2}$ is normal. Todorčević proved if X is an entangled set, then all finite power of $X[\leq]$ are hereditarily Lindelöf, hence $(X[\leq])^{2}$ is normal. We construct from CH a subset X of reals, such that X is not a $\lambda$-set (hence not a Q-set), X is not 2-entangled, but still $(X[\leq])^{2}$ is normal. We will also prove the natural analogue of Przymusiński’s theorem for $\lambda$-sets, namely, if X is a $\lambda$-set then $(X[\leq])^{2}$ is pseudo-normal.
@@
$Q$-set, $\lambda$-set, entangled sets, Sorgenfrey topology, normality
@@
Primary: 54D15, 54G20 Secondary: 03E50, 03E05, 03E75
The first author acknowledges support from NSERC Grant RGPIN-2025-06324.
@@ 1. Introduction
A $Q$-set is an uncountable subset of the reals with the property that every subset is $G_{\delta}$ (equivalently, $F_{\sigma}$) in the relative Euclidean subspace topology. The notion was first introduced in [4] and much later $Q$-sets were used to produce consistent counterexamples to the Normal Moore Space Conjecture [2]. For example, if $X\subseteq 2^{\omega}$, the set of branches $b_{x}=\{x\upharpoonright n:n<\omega\}$ in $2^{<\omega}$ is an almost disjoint family, and the corresponding $\Psi$-space on $2^{<\omega}$ is normal if and only if $X$ is a $Q$-set. Similarly, if $X$ is a subset of the $x$-axis in the upper half plane of ${\mathbb{R}}^{2}$, then the Moore-Niemyitzki Plane topology (a.k.a. the bubble space) is normal if and only if $X$ is a Q-set. For more on these examples see [7]. On the other hand Prymusiński [6] proved that if $X$ is a Q-set endowed with the Sorgenfrey Topology then $X^{2}$ is normal.
@@
We first observe that the converse to Przymusiński’s theorem does not hold: Todorčević observed that if $X$ is an entangled set [8] of reals and is endowed with the Sorgenfrey topology, then every finite power is both hereditarily Lindelöf and hereditarily separable [9]. Recall that CH implies both the existence of entangled sets of reals and the non-existence of $Q$-sets. So $(X[\leq])^{2}$ may be normal even if $X$ is not a $Q$-set.
There is the weaker notion of a set of reals $X$ being a $\lambda$-set [5]. $X$ is a $\lambda$-set if every countable subset is relatively $G_{\delta}$ in the Euclidean topology. While the existence of a $Q$-set of reals is independent of ZFC, there are $\lambda$-sets in ZFC. Analogous to the $Q$-set results mentioned above, the $\Psi$-space over the set of branches of a subset $X\subseteq 2^{\omega}$ is pseudo-normal (meaning that every pair of closed sets, one of which is countable, can be separated by disjoint open sets) if and only if $X$ is a $\lambda$-set and also $X$ is a $\lambda$-set if and only if the Moore-Niemytzki plane over $X$ is pseudo-normal (see [10]).
We provide some partial converses to Przymusiński’s theorem underscoring necessary conditions for normality of $(X[\leq])^{2}$. We show that the expected analogous result for the Sorgenfrey topology on a $\lambda$-set holds. Namely, if $X$ is a $\lambda$-set of reals, then $(X[\leq])^{2}$ is pseudo-normal. We also construct, assuming CH, a set of reals for which $X$ is concentrated on a countable dense subset (so not a $\lambda$-set) and so that $X$ is also not 2-entangled, but still $X[\leq]^{2}$ is normal.
@@
Let $X$ be a topological space. Let $A,B$ be subsets in $X$, AA has the covering property with respect to $B$ iff there exists a countable open cover $\{U_{n}\}_{n<\omega}$ such that $\overline{U_{n}}\cap B=\emptyset$. If there is no ambiguity in the context, we just say $A$ has the covering property.
@@
Since we will have two topologies on $\mathbb{R}$, we will use the following notation for closure and compliments.
Definition 1.3.
For $X\subseteq\mathbb{R}$ be a subset of reals, we denote $\overline{X}^{E}$ as the Euclidean closure and $\operatorname{int}_{E}(X)$ as Euclidean interior. We denote $\overline{X}^{\sigma}$ and $\operatorname{int}_{\sigma}(X)$ to denote the Sorgenfrey closure and Sorgenfrey interior respectively.
The Sorgenfrey topology is finer than the Euclidean topology, i.e. every Euclidean open set is also Sorgenfrey open but not vice versa.
@@
All of our topological terminology is standard and we refer the reader to Engelking [3] for any undefined notions.
@@ 2. Q-sets and the Sorgenfrey square
Let $A,B$ be subsets of reals. For any monotone decreasing function $f:A\to B$, there is a countable set $B_{0}\subseteq B$ such that $f\upharpoonright(A\setminus f^{-1}(B_{0})):A\setminus f^{-1}(B_{0})\to f(A)\setminus B_{0}$ is a Euclidean homeomorphism.
If, further, $f$ is strictly decreasing, then there is a countable $A_{0}\subseteq A$ such that $f\upharpoonright(A\setminus A_{0}):A\setminus A_{0}\rightarrow f(A\setminus A_{0})$ is a Euclidean homeomorphism.
@@
Suppose $X,Y$ are sets of reals. Suppose that
@@
    there exists a strictly decreasing Euclidean homeomorphism $f:X\to Y$.
@@
If there exists $\mathcal{U},\mathcal{V}\subseteq X\times Y$ Sorgenfrey open such that $\mathcal{U}\supseteq G_{f|_{A}},\mathcal{V}\supseteq G_{f_{B}}$ and $\mathcal{U}\cap\mathcal{V}=\emptyset$, then $A$ is relatively $F_{\sigma}$ in $X$ in the Euclidean topology.
@@
Note that $A=\bigcup_{m<\omega}A_{m}$ where $A_{m}=\{a\in A:U((a,f(a)),1/m)\subseteq\mathcal{U}\}$, we claim that
@@
$\overline{A_{m}}^{E}\cap X\subseteq A$
With this claim, we have $A=\bigcup_{m<\omega}(\overline{A_{m}}^{E}\cap X)$ which is a relatively Euclidean $F_{\sigma}$ set. So it suffices to prove the claim.
Proof of 2.3.
Assume otherwise, if $a_{n}\to b$ for $a_{n}\in A_{m}$ and $b\in B$. There exists an $r>0$ such that $U((b,f(b)),r)\subseteq\mathcal{V}$. By continuity of $f$, we have $f(a_{n})\to f(b)$ as well. Let $\epsilon<\min(1/m,r)$, without loss of generality there exists $a_{n_{0}}$ such that $b>a_{n_{0}}>b-\epsilon$ and $f(b)<f(a_{n_{0}})<f(b)+\epsilon$. Since $\epsilon<\min(1/m,r)$, we have
$f(b)<\mathbf{f(a_{n_{0}})}<f(b)+\epsilon<f(b)+r$ $\implies\mathbf{f(a_{n_{0}})}\in[f(b),f(b)+r)$
$a_{n_{0}}<\mathbf{b}<a_{n_{0}}+\epsilon<a_{n_{0}}+1/m$ $\implies\mathbf{b}\in[a_{n_{0}},a_{n_{0}}+1/m)$
Where the first expression implies $(b,f(a_{n_{0}}))\in U((b,f(b)),r))\subseteq\mathcal{V}$ However, the second expression implies $(a_{n_{0}},f(a_{n_{0}}))$ is also a point in $U((a_{n_{0}},f(a_{n_{0}})),1/m)\subseteq\mathcal{U}$, but $\mathcal{U}$ and $\mathcal{V}$ should have empty intersection hence this is a contradiction. ∎
@@
Let $(X[\leq])^{2}$ be normal. If there exists a strictly decreasing Euclidean homeomorphism $f:X\to X$, then $X$ is a $Q$-set.
@@
Since $f$ is an Euclidean homeomorphism, its graph $G_{f}$ is Euclidean closed and so Sorgenfrey closed. Since $f$ is strictly decreasing, $G_{f}$ is also Sorgenfrey discrete. Hence any subset of $G_{f}$ is Sorgenfrey closed. In particular, for any $A\subseteq X$, $G_{f|_{A}}$ and $G_{f|_{X\setminus A}}$ can be separated by Sorgenfrey open sets (by our normality assumption). By 2.2, $A$ is relatively Euclidean $F_{\sigma}$ in $X$. Hence $X$ is a $Q$-set. ∎
@@
By 2.1, there exists a countable set $C$ such that $f|_{X\setminus C}:X\setminus C\to f(X\setminus C)$ is a Euclidean homeomorphism. Then let $g=f|_{X\setminus C}$, the graph $G_{g}$ of $g$ is closed and discrete. Since $(X[\leq])^{2}$ is normal, any two subsets of $G_{g}$ can be separated by Sorgenfrey open subsets of $X^{2}$. In particular, let $X^{\prime}=X\setminus C$, take any subset $A\subseteq X^{\prime}$, $G_{g|_{A}}$ and $G_{g|_{X\prime\setminus A}}$ can be separated by open sets $\mathcal{U},\mathcal{V}$. Take their intersection to $(X^{\prime})^{2}$, $\mathcal{U}\cap(X^{\prime})^{2}$ and $\mathcal{V}\cap(X^{\prime})^{2}$ are still Sorgenfrey open in $(X^{\prime})^{2}$. By 2.2, $A$ is relatively Euclidean $F_{\sigma}$ in $X^{\prime}$, hence $X^{\prime}$ is a $Q$ set as required. ∎
Baumgartner proved that PFA implies that any two $\aleph_{1}$-dense sets of reals are order isomorphic [1]. This consequence of PFA has come to be referred to as Baumgartner’s Axiom (BA). From the proof of the above proposition we obtain the following corollary:
@@
Any uncountable $X$ contains an $\aleph_{1}$-dense subset $X^{\prime}$. By BA it is order isomorphic to its reverse order. So by the proof of Proposition 2.5 there is a countable $C$ such that $X^{\prime}\setminus C$ must be a $Q$-set.∎
@@
Let $f:A\to\mathbb{R}$ be strictly decreasing on uncountable $A\subseteq\mathbb{R}$, then there is a strictly decreasing homeomorphism $F$ on $\overline{A}^{E}\setminus C$ for some countable set $C$ such that $f(a)=F(a)$ for all $a\in A\setminus C$.
@@
$f$ can be extended to a decreasing function on $\overline{A}^{E}$. Note that this function is not necessarily strictly decreasing. By 2.1, we can remove a countable set $B_{0}$ such that $f$ is a Euclidean homeomorphism $F$ on $\overline{A}^{E}\setminus f^{-1}(B_{0})$ hence strictly decreasing. We will show $f^{-1}(B_{0})$ is countable by showing for every $b\in B_{0}$, $f^{-1}(b)$ has at most 3 points. Assume there are 4 points $x<y<z<w$ in $f^{-1}(b)$. Note that between $x$ and $w$ there is at most 1 point in $A$, since for every $z_{1}<z_{2}\in A$,
@@
but $f$ is strictly decreasing on $A$, so there can only be at most 1 point in $A$. Hence, there are no points in $\overline{A}^{E}\setminus A$. But $y,z$ can not be both in $A$, we have a contradiction. Let $C=f^{-1}(B_{0})$, we have the result. ∎
We also recall definition of entangled sets from [8]. Let $L\subseteq\mathbb{R}$, denote $(L)^{n}\subseteq L^{n}$ as the set of all $\langle x_{0},\dots,x_{n-1}\rangle\in L^{n}$, such that $x_{0}<x_{1}<\dots<x_{n}$. A set $L\subseteq\mathbb{R}$ is $\mathbf{n}$-entangled iff for any set of disjoint $n$-tuples $A\subseteq(L)^{n}$ of size $\omega_{1}$ and for every $s\in 2^{n}$ there exists $x,y\in A$ such that $\forall i\leq n,(x_{i}<y_{i})\iff s_{i}=0$. $L$ is entangled iff $L$ is $\mathbf{n}$-entangled for all $n<\omega$. Let $s\in 2^{n}$, $A\subseteq(L)^{n}$ is said to have type $s$ iff $\forall i\leq n(x_{i}<y_{i})\iff s_{i}=0$.
The following lemma is proven in [8] (see property (5) in the proof of Theorem 1 in [8]).
@@
[8] Let $E\subseteq\mathbb{R}$ be such that for any continuous function $f$ from a $G_{\delta}$ subset of ${\mathbb{R}}^{n}$ into $\mathbb{R}$}, $(n<\omega)$ there is an $\alpha<2^{\omega}$ such that
$\forall\beta\geq\alpha,f(\{r_{\gamma}:\gamma<\beta\}^{n})\cap E\subseteq\{r_{\gamma}:\gamma<\beta\}$
@@
    For any continuous function $f$ from a $G_{\delta}$ subset of ${\mathbb{R}}^{n}$ into $E$, $(n<\omega)$ there is an $\alpha<2^{\omega}$ such that
    $\forall\beta\geq\alpha,f(\{r_{\gamma}:\gamma<\beta\}^{n})\cap E\subseteq\{r_{\gamma}:\gamma<\beta\}$
@@
    $E$ contains $\mathbb{Q}$ and is concentrated on $\mathbb{Q}$}.
Recall that $E$ is concentrated on $\mathbb{Q}$ means that $E\setminus U$ is countable for every open set $U\supseteq\mathbb{Q}$, so $E$ is not a $\lambda$-set since $\mathbb{Q}$ is not a relative $G_{\delta}$-set in $E$.
@@
$\mathcal{F}=\{$ $f:U\to\mathbb{R}:U\subseteq\mathbb{R}^{n},n<\omega,U\text{ is }G_{\delta}\}\}$
@@
For every $\alpha<\omega_{1}$ let $n_{\alpha}$ be the parity of the function $f_{\alpha}$ and define the following set:
$C_{\alpha}$ $=\bigcap_{\beta\leq\alpha}A_{\beta}\setminus\mathbb{Q}.$
Note that $C_{\alpha}$ is uncountable by the Baire Category Theorem. We now define elements $x_{\alpha}\in\mathbb{R}$ and countable sets $I_{\alpha}\subseteq{\mathbb{R}}$ recursively as follows:
Stage 0: Let $I_{0}=f_{0}({\mathbb{Q}}^{n_{0}})$ and take $x_{0}\in C_{0}\setminus I_{0}$.
Stage $\bm{\alpha<\omega_{1}}$: We have $x_{\beta}$ and $I_{\beta}$ defined already for all $\beta<\alpha$. For any $\beta\leq\omega_{1}$, let $n_{\beta}$ be the parity of $f_{\beta}$, and let
$X_{\alpha}$ $={\mathbb{Q}}\cup\{x_{\beta}:\beta<\alpha\}$
$I_{\alpha}$ $=\bigcup_{\beta\leq\alpha}f_{\beta}(X_{\alpha}^{n_{\beta}})$
Note that all of the above sets are countable, and $C_{\alpha}$ is uncountable. Hence, we can choose $x_{\alpha}\in C_{\alpha}\setminus(X_{\alpha}\cup I_{\alpha})$.
@@
Verify Property (1): Fix $n<\omega$, for any continuous function $g:A\to E$, by choice of $x_{\alpha}$, we have that for any $\beta\geq\alpha$, $f_{\alpha_{0}}(\{x_{\gamma}:\gamma<\alpha\}^{n})$ can not take on any values in $\{x_{\beta}:\beta\geq\alpha_{0}\}$. Indeed for all $\beta>\alpha_{0},$ $x_{\beta}\not\in I_{\alpha}\supseteq f_{\alpha_{0}}((\{x_{\alpha}:\alpha<\alpha_{0}\}\cup\mathbb{Q})^{n})$
Verify Property (2): That $E$ is concentrated on $\mathbb{Q}$ follows easily from the fact that for every for every open set containing $\mathbb{Q}$} is enumerated as $A_{\alpha}$ and so $C_{\beta}\subseteq A_{\alpha}$ for all $\beta\geq\alpha$.
As mentioned above, the set $E$ constructed is not a $\lambda$-set so also not a $Q$-set.
By 2.8, the set $E$ is also entangled and in particular $2$-entangled. By 2.7 this implies no monotone functions can be defined on an uncountable subset of $E$.
For our next observation, we will use the following result, proven by Przymusiński (see the proof of Theorem 1 in [6]).
Lemma 2.10.
Let $X\subseteq\mathbb{R}$ be any subset of reals. Fix any Sorgenfrey closed set $A,B\subseteq X^{2}$, and a Sorgenfrey open cover $\mathcal{A}$ of $A$. For any $z\in B^{c}$ choose $n_{z}\in\mathbb{N}$ such that $U(z,1/n_{z})\cap B=\emptyset$, this is possible since $B$ is Sorgenfrey closed. Let $A\cap\overline{B}^{E}=\bigcup_{q\in\mathbb{Q}}A_{q}$ where
$A_{q}=\{z=(x,y)\in A_{1}:x+y<q<x+y+1/n_{z}\}$
For any $q\in\mathbb{Q}$, we have
  1. (1)
    $A_{q}=G_{q}\cup L_{q}$, where $L_{q}$ can be covered by countably many Sorgenfrey open sets of $\mathcal{A}$ and $G_{q}$ is a graph of some strictly decreasing function,
  2. (2)
    If every $G_{q}$ can be covered by countably many Sorgenfrey open sets in $\mathcal{A}$, then $X[\leq]^{2}$ is normal.
Recall, Todorčević noted [9] that if $E$ is entangled, then $E[\leq]$ is hereditarily Lindelöf in all finite powers. In relation to this we have: